OpenStudy (anonymous):

the difference of two ages of two boys x and y is 3 years. And the sum of their ages is 27. Find their ages. please help me ..

OpenStudy (anonymous):

y = 12 x = 15

OpenStudy (anonymous):

but how did you get it ? may i know?

OpenStudy (anonymous):

Of course! (: You set up two equations.

OpenStudy (anonymous):

1. x + y = 27 2. x - y = 3

OpenStudy (anonymous):

These equations efficiently put all the information into a usable format.

OpenStudy (anonymous):

With this system of equations, you'd make another equation in the perspective of y.

OpenStudy (anonymous):

To do this, you get your first equation (x + y = 27) and subtract y so you get x= 27 - y.

OpenStudy (anonymous):

i really hate math ! T.T

OpenStudy (anonymous):

Haha well you're in luck! It's my favorite subject (:

OpenStudy (anonymous):

Anyways with your new equation (x = 27 - y) you substitute the x in your second equation (x - y = 3 You end up with (27 - y) - y = 3

OpenStudy (anonymous):

Good so far?

OpenStudy (anonymous):

i got it . thank you so much sis .. :)) Godbless

OpenStudy (anonymous):

Aw no problem! All you have to do from this point on is combine the y's, solve for y, and then subtract your answer for y from 27 to get x. HAVE A LOVELY DAY!

OpenStudy (anonymous):

but sis about this: A man is 4 times as old as his daughter. four years from now, the sum of their age will be 78. find their present ages.

OpenStudy (anonymous):

Alright give me a couple minutes to plan it out on paper

OpenStudy (anonymous):

Okay got it. What I started out with was m = 4d (m for man and d for daughter).

OpenStudy (anonymous):

What you need to understand is that this^ equation is in the PRESENT.

OpenStudy (anonymous):

To find the ages you need to use another equation which is 4 YEARS from now. To do this, you can use this equation. (4d + 4) + (d + 4) = 78

OpenStudy (anonymous):

Basically what you do is combine like terms and get 5d + 8 + 78. Then subtract 8 and divide 5 and you see that the daughters age is 14. Just multiply 14 by 4 to get the age of her dad and you're done!

OpenStudy (anonymous):

5d + 8 = 78* lol woops

OpenStudy (anonymous):

thank you so much ! .. I pray for More knowledge and wisdom of you.

OpenStudy (anonymous):

No problem love (:

OpenStudy (anonymous):

this one, my last question, the sum of 3 consecutive odd integers is 27. what are the numbers?

OpenStudy (anonymous):

haha my favorite. Your equation is n + n + 2 + n + 4 = 27

OpenStudy (anonymous):

combine like terms and you get 3n + 6 = 27. subtract 6 3n = 21. divide by 3 n = 7

OpenStudy (anonymous):

your numbers are 7, 9 and 11

OpenStudy (anonymous):

but why n+4?

OpenStudy (anonymous):

Okay. Since they are consecutive odd numbers, the easiest way to describe them in an equation in using a 'like term' if that makes any sence.

OpenStudy (anonymous):

For example. to decribe 1 + 3 + 5 = 9 using variables, you would do n + n + 2 + n + 4 Of course using the value of n as the smallest number.

OpenStudy (anonymous):

ah. now i know. thanks a lot sis, may God Bless you for helping me :))

OpenStudy (anonymous):

Don't mention it (:

OpenStudy (anonymous):

bye sis :))

OpenStudy (anonymous):

bye c: